Tim Howard has been offered the opportunity of investing $37,263 now. The investment will earn 7% per year and at the end of that time will return Tim $89,800.

How many years must Tim wait to receive $89,800?
Answer #1

Given,

Present value (PV) = 37,263

Rate of interest (r) = 0.07

Future value (FV) = 89,800

Number of years (n) =?

Now by the formula as below:

FV = PV (1 + r) ^ n

89,800 = 37,263 (1 + 0.07) ^ n

(89,800 / 37,263) = 1.07 ^ n

2.409897 = 1.07 ^ n

n = [In (2.409897) / In (1.07)]

   = 0.879584008 / 0.06765864847

   = 13.000

Answer: 13 years